# Optimisation models for the design of multiple self-consumption loops in semi-rural areas

### v2

@@ 4 Application of the models
This section details the data and assumptions used to set up realistic instances and their corresponding consumption and production profiles. The method to generate these instances is designed to mimic real-case scenarios. These instances are then reused in a business-oriented test protocol, which is described in Section 4.2.
@@ 4.1.1 Consumption load
To generate a group of consumers, we divided this group into two consumption sub-groups, namely ”Household” and ”Pro: Professional.” These categories echo the separation made by the Enedis Open Data platform, which provides an average profile of each category. The distribution of consumers in each category is manually determined based on the scenario chosen for simulation (see Section 4.2). All consumers in each category are then provided with a consumption profile that is normally distributed around the Enedis reference profile load corresponding to that category.
@@ 4.3 Evaluation metrics
Each model presented in Section 3 is applied to generated instances. Aside from objective value and solving time, results are analysed through several evaluation metrics defined below.
@@ 4.3.4 Average number of members and installed power per loop
$N_{\text{mean}}=\frac{1}{N_{l}}\sum_{h\in\mathcal{H}}\sum_{i\in[n]}v_{h}a_{i}^{h}.\$ (14)
The average number of loops (14) can be put into perspective with the number of actors that are not participating in any loop, telling how created loops include actors. Another indicator underlines how complete loops are in terms of installed power. This indicator should always be analysed in relation to the legal installed power limit set initially to solve the problem (15):
$P_{\text{mean}}=\frac{1}{N_{l}}\sum_{h\in\mathcal{H}}\sum_{i\in[n]}v_{h}a_{i}^{h}P_{i}^{\text{inst}}.$ (15)
@@ 4.3.5 Total benefit realised through self-consumption
This indicator is obtained through the calculation of realised benefits by each prosumer who is involved in a self-consumption loop relative to a zero loop configuration (16)
@@ 5 Results and discussion
5.1 Scalability on the temporal dimension
(a) Along time horizon (15 prosumers)
(b) Along number of prosumers (7 days)
@@ 5.1 Scalability on the temporal dimension
As explained in the model presentation, extending the time horizon significantly increases the complexity of the problem, particularly because each time step can influence whether or not an actor should be included in a collective self-consumption operation. Whether an actor belongs to or is excluded from a self-consumption operation is determined for the entire optimisation time horizon and cannot be changed from one time step to another. The existence of associated integer variables links each time step in the optimisation problem, and in the case of the compact formulation, requires all time steps to be solved within the same problem. The Benders extended reformulation allows for the extraction of integer variables and fixing them in the master problem, thereby enabling each time step of the sub-problem to be easily solvable independently. The increase in complexity over time horizon is the reason why a Benders decomposition of the model has been proposed. Figure 2(a) illustrates the evolution of the resolution time of the single-loop models for different time horizons, comparing the compact formulation with the extended formulation.
@@ 5.2 Solution sensitivity analysis
In addition to runtime performance, another important focus is on the optimal solutions derived from the multi-loop model from a business perspective. This entails evaluating the solutions in terms of their cost-effectiveness and assessing the influence of business-related parameters on proposed solutions.
5.2.1 Influence of legal maximum installed power $P_{\mathrm{leg}}$
@@
Figure 4: Influence of the legal power on businesses indicators
Figure 3(a) and 3(c) respectively illustrate the impact of this parameter on the average size of the corresponding loops (in terms of installed power) and the benefit of the formed collective self-consumption loops.
@@
Figure 5: Influence of distribution density on businesses indicators
@@ 5.2.2 Influence of actor density
Increasing actors’ density favours the formation of larger collective self-consumption communities. Figure 4(a) illustrates a slight increase in the installed power of loops as clusters are brought closer together. This reflects the fact that separated clusters have a greater chance of being at a reasonable distance in denser configurations. This observation is confirmed by analyzing the average number of actors per loop in Figure 4(b). Since larger loops ensure a greater amount of locally produced electricity, the overall gain realised is also directly affected by an increase in density, as shown in Figure 4(c).
@@ 5.3 Influence of exposition and period of the year
(a) Self-production rate
(b) Self-consumption rate
@@
Figure 6: Influence of exposition on businesses indicators
Solar panel exposure directly impacts production volumes as it influences the production yield of plants. Even though the best configurations logically favour better economic gain (see Figure 5(c)), dealing with different expositions may result in changes in business indicators. In the worst exposition case, the total production amount is lower than consumption needs, whereas it is higher with the best exposition configuration. This results in an inversion in terms of self-production rate and self-consumption rate, as Figure 5(a) and 5(b) demonstrate.
@@
Figure 7: Influence of the period of the year on business indicators
The previously mentioned inversion is not observed when the changes in production come from considering two different periods of the year (see Figure 6(a) and 6(b)). This can be explained by the fact that the period of the year has less impact on production than the panel exposition. When production can be divided by 10 due to poor exposition, the ratio between summer and winter is less than 2. In the presented instances, this means that in both summer and winter periods, production volume exceeds consumption volume. Still, considering the summer period seems to be more in favour of self-production than self-consumption.
@@ 6 Conclusion and future work
Overall, the article, powered by presented contributions, provides comprehensive insights into the optimisation of collective self-consumption communities, considering both technical and business aspects. We conducted a performance-oriented analysis on instances of various sizes and also performed a sensitivity analysis on business-related parameters, showing how these models can support decision-making and help understand solar plant installation contexts’ influence on self-consumption loop design.
@@ 7 Acknowledgment
8 Funding
This work was supported by the French Ministry of Economy, Finance and Industrial and Digital Sovereignty through a France Relance project.

### v3

@@ 4 Application of the models
This section details the data and assumptions used to set up realistic instances and their corresponding consumption and production profiles. The method to generate these instances is designed to mimic real-case scenarios. These instances are then reused in a test protocol on operational-specific metrics, which is described in Section 4.2.
@@ 4.1.1 Consumption load
To generate a group of consumers, we divided this group into two consumption sub-groups, namely ”Household” and ”Pro: Professional.” These categories echo the separation made by the Enedis Open Data platform, which provides an average profile of each category. The distribution of consumers in each category is manually determined based on the scenario chosen for simulation (see Section 4.2). All consumers in each category are then provided with a consumption profile that is normally distributed around the Enedis reference profile load corresponding to that category. To do so, a rolling window mean using a 3 steps window size is applied. Some variations are added to this averaged consumption using a uniform random variable and a variation factor of 30%.
@@ 4.2 Test protocol
Instances involving 10 actors and having the same configuration as the reference instance presented in 5 are also solved for different time horizons to evaluate how the models behave in solving larger problems. With the same objective, instances having several number of prosumers and a time horizon of 7 days are solved.
Table 7 indicates the different sizes - in terms of number of variables and constraints for all instances averaged on 100 replicates for each of them. All Gurobi generated instances are available as .lp files at Chasseray et al., (2025).
Table 7: Instances sizes (average computed on 100 replicates for each configuration and model)
Constraints (×$\times 10^{3}$) Variables (×$\times 10^{3}$) Integer variables Configuration MLCol MLCpct SLCpct MLCol MLCpct SLCpct MLCol MLCpct SLCpct reference 332 470 243 111 201 121 1198 6714 735 dens_0_1 154 445 184 80 173 92 375 6529 735 dens_2 574 521 335 156 249 167 2168 7367 735 period_sum 441 475 243 111 203 121 1198 6856 735 power_1000 127 475 243 110 203 121 831 6856 735 power_5000 431 475 243 111 203 121 1541 6856 735 config_wc 226 475 243 110 203 121 1962 6856 735 10_act_1_d 5 8 5 2 4 2 386 500 34 10_act_10_d 45 73 47 21 34 23 386 1580 250 10_act_14_d 63 102 66 29 48 33 386 2060 346 10_act_30_d 138 218 140 62 102 70 386 3980 730 10_act_60_d 284 434 281 124 204 140 386 7580 1450 10_act_90_d 442 651 421 186 305 210 386 11180 2170 10_act_180_d - 1301 843 - 610 419 - 21980 4330 10_act_270_d - 1952 1264 - 914 629 - 32780 6490 10_act_360_d - 2602 1686 - 1219 839 - 43580 8650 10_act_7_d 31 51 33 15 24 16 386 1220 178 15_act_7_d 38 106 44 20 41 22 421 2137 183 20_act_7_d 39 166 54 24 60 27 339 2860 188 30_act_7_d 56 354 80 35 115 40 462 5165 198
@@ 4.3 Evaluation metrics
Each model presented in Section 3 is applied to generated instances. Aside from objective value and solving time, results are analysed through several evaluation metrics defined below. These KPI are only used for a post optimisation analysis of the loops, as benefit balance within the loops is discussed but left for future work.
@@ 4.3.4 Average number of members and installed power per loop
The average number of loops
$N_{\text{mean}}=\frac{1}{N_{l}}\sum_{h\in\mathcal{H}}\sum_{i\in[n]}v_{h}a_{i}^{h}$
can be put into perspective with the number of actors that are not participating in any loop, indicating how created loops include actors. Another indicator is the average installed power per loop, which underlines how complete loops are in terms of installed power:
$P_{\text{mean}}=\frac{1}{N_{l}}\sum_{h\in\mathcal{H}}\sum_{i\in[n]}v_{h}a_{i}^{h}P_{i}^{\text{inst}}.$
This indicator should always be analysed in relation to the legal installed power limit set initially to solve the problem.
@@ 4.3.5 Total benefit realised through self-consumption
This indicator is obtained from realised benefits by each prosumer who is involved in a self-consumption loop relative to a zero loop configuration:
@@ 5 Results and discussion
5.1 Scalability on the temporal and geographical dimensions
(a) Across time horizon (10 actors)
(b) Across number of actors (7 days)
@@ 5.1 Scalability on the temporal and geographical dimensions
As explained in the model presentation, extending the time horizon significantly increases the complexity of the problem, since each time step can influence whether or not an actor should be included in a collective self-consumption operation. Whether an actor belongs to or is excluded from a self-consumption operation is determined for the entire optimisation time horizon and cannot be changed from one time step to another. The existence of associated integer variables links each time step in the optimisation problem, and in the case of the compact formulation, requires all time steps to be solved within the same problem. The Benders extended reformulation allows for the extraction of integer variables and fixing them in the master problem, thereby enabling each time step of the sub-problem to be easily solvable independently. The increase in complexity over time horizon is the reason why a Benders decomposition of the model has been proposed. Figure 3(a) illustrates the evolution of the resolution time of the single-loop models for different time horizons, comparing the compact formulation with the extended formulation.
@@
5.2 Root-node relaxation tightness
(a) Across time horizon (10 actors) - SLCpct model - Three outliers removed for 10 days (value: 7.46), 14 days (value: 21.53) and 30 days (value: 21.50) configurations
(b) Across number of actors (7 days) - SLCpct model - One outlier removed for 10 actors configuration (value: 11.20)
(c) Across time horizon (10 actors) - MLCpct (left) and MLCol (right) models. One outlier removed for 30 days configuration (MLCol - value: 17.10)
(d) Across number of actors (7 days) - MLCpct (left) and MLCol (right) models.
Two outliers removed at 10 actors (MLCol - value: 10.27) and 30 actors configuration (MLCpct - value: 3.64)
Figure 4: Root gap distribution by solving method across configurations (some outliers have been removed from the plots for clarity purpose)
We analyse in this section the quality of the formulation through the root-node gap which provides a measure of how tight the different relaxations are. Figure 4 illustrates how the root-node gap evolves with respect to the time horizon and the number of considered prosumers for the different problem formulations. Focusing first on single-loop models, Figure 4(b) shows that increasing the number of prosumers results in smaller relative gaps, thus with linear relaxations that are tighter around integer solutions. Interestingly, as illustrated in Figure 4(a), the root-node gap is maximal for ten prosumers between 14 and 30 days of horizing, with shorter and longer horizons both resulting in smaller gaps and tight formulations, which leads us to conjecture that longer planning horizons increases the contrast in the economic interest to select actors, reducing the root-node fractionality.
Having a lower relative root gap however, does not imply that the problem can be solved in a shorter time, as shows Figure 3(a) for the single loop formulation. Nevertheless, it indicates that the resolution should be easier relatively to the size of the mixed-integer formulation (number of variables and constraints).
This tendency does not hold for multiple-loop compact formulations. In these cases, the root gap tends to increase with longer time horizons (greater than 30 days). The same effect is particularly true when increasing the number of prosumers, especially from 20 to 30 prosumers, where a significant root gap difference is observed.
What remains noteworthy, though, is that the reformulated problem using clique enumeration exhibits the opposite behavior. Similar to single-loop models, it exhibits a peak root gap around 14 days of time horizon and decreases for larger time horizons.
Figure 4(c) and Figure 4(d) highlight the benefits of the clique generation formulation. This approach results in lower root gaps than the compact model once the number of prosumers becomes large (i.e., greater than 20). Furthermore, unlike the compact formulation, clique generation not only produces smaller root gaps but also tends to further reduce them as either the time horizon or the number of prosumers increases, hinting that our approaches will produce high-quality solutions and guarantees even without running a full branch-and-cut algorithm.
@@ 5.3 Solution sensitivity analysis
In addition to runtime performance, another important focus is on the solutions derived from the multi-loop model from the application perspective. This entails evaluating the solutions in terms of their cost effectiveness and assessing the influence of operational parameters on proposed solutions.
5.3.1 Influence of the legal maximum installed power
@@
Figure 5: Influence of the legal power on operational indicators
Figure 5(a) and 5(c) respectively illustrate the impact of this parameter on the average size of the corresponding loops (in terms of installed power) and the benefit of the formed collective self-consumption loops.
@@
Figure 6: Influence of distribution density on operational indicators
@@ 5.3.2 Influence of actor density
Increasing the actor density favours the formation of larger collective self-consumption communities. Figure 6(a) illustrates a slight increase in the installed power of loops as clusters are brought closer together. This reflects the fact that separated clusters have a greater chance of being at a reasonable distance in denser configurations. This observation is confirmed by analyzing the average number of actors per loop in Figure 6(b). Since larger loops ensure a greater amount of locally produced electricity, the overall gain realised is also directly affected by an increase in density, as shown in Figure 6(c).
@@
(a) Root gap evolution along actors density for SLCpct model
(b) Root gap evolution along actors density for MLCpct (left) and MLCol (right) models
Figure 7: Root gap distribution by solving method along actors density
Figure 7 shows the relative root gap for group of instances having different prosumers density. We observe that the compact formulation performs better for denser configurations, as having close actors is similar to relaxing the distance constraint, leading to an actors selection problem driven only by the benefit of loops and constrained by power constraint. This figure also shows that the formulation involving clique generation (MLCol) is not suited for denser problems, as having closer actors conducts to a high number of possible cliques, and a bigger number of variables. For sparse configurations however the clique generation performs better than the compact formulation, as many invalid actors combinations are not evaluated.
@@ 5.4 Influence of exposition and period of the year
(a) Self-consumption rate
(b) Covering rate
@@
Figure 8: Influence of exposition on operational indicators
Solar panel exposure directly impacts production volumes as it influences the production yield of plants. Even though the best configurations logically favour better economic gain (see Figure 8(c)), dealing with different expositions may result in changes in operational indicators. In the worst exposition case, the total production amount is lower than consumption needs, whereas it is higher with the best exposition configuration. This results in an inversion in terms of self-production rate and self-consumption rate, as Figure 8(a) and 8(b) demonstrate.
@@
Figure 9: Influence of the period of the year on operational indicators
The previously mentioned inversion is not observed when the changes in production come from considering two different periods of the year (see Figure 9(a) and 9(b)). This can be explained by the fact that the period of the year has less impact on production than the panel exposition. When production can be divided by 10 due to poor exposition, the ratio between summer and winter is less than 2. In the presented instances, this means that in both summer and winter periods, production volume exceeds consumption volume. Still, considering the summer period seems to be more in favour of self-production than self-consumption.
@@ 6 Conclusion and future work
Overall, the article, powered by presented contributions, provides comprehensive insights into the optimisation of collective self-consumption communities, considering both technical and operational aspects. We conducted a performance-oriented analysis on instances of various sizes and also performed a sensitivity analysis on operational parameters, showing how these models can support decision-making and help understand solar plant installation contexts’ influence on self-consumption loop design.
@@ 7 Acknowledgment
8 Compliance with Ethical Standards
8.1 Funding
This study was founded by the French Ministry of Economy, Finance and Industrial and Digital Sovereignty through a France Relance project.
8.2 Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.